OpenStudy (anonymous):

The basketball team scored 12 points the first game and 47 points the eighth game of the season. How many points will they earn with this pattern on the 14th game of the season?

OpenStudy (anonymous):

I'm going to guess that this is an arithmetic sequence, geometric would be huge. 47 - 12 = 35..............(35/7) = 5 five more games is 25 more points than 47. 47 + 25 = 72 <--------------

OpenStudy (anonymous):

The options are 68,71,74, and 77

OpenStudy (anonymous):

= 77

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

You can refer @ZhangYan's solution!

OpenStudy (anonymous):

but he didnt get the same answer?

OpenStudy (anonymous):

Just wrong calculation!

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

ZhangYan he go the answer of of an other person he didnt even do calculations